What value of the variable x us associated with:
a. Y=0
b. y=1
c. Y=-8
the equation is y=x cubed

Answer by stanbon(57387)

What value of the variable x us associated with:
a. Y=0
Solve x^3 = 0
Ans: x = 0
-----------------
b. y=1
Solve x^3 = 1
x = 1
-------------------
c. Y=-8
Solve x^3 = -8
x = -2
===================================
the equation is y=x cubed
Cheers,
Stan H.